Part 1 Suppose a graduate student receives a non-subsidized student loan of $13,000 for each of...

Part 1

Suppose a graduate student receives a non-subsidized student loan of $13,000 for each of the 4 years the student pursues a PhD. If the annual interest rate is 5% and the student has a 10-year repayment program, what are the student's monthly payments on the loans after graduation? (Round your answer to the nearest cent.)

$____________

Part 2

Samuel Ng received a 3-year subsidized student loan of $12,000 at an annual interest rate of 5%. What are Samuel's monthly loan payments for this loan when he graduates from college in 2 years? (Round your answer to the nearest cent.)

$_____________
Answer #1

Part 1

We will use the excel's PMT, PV and FV function to solve the problem

The student receives a non-subsidized student loan of $13,000 for each of the 4 years

Annual Interest rate = 5%

Using excel's FV function to arrive at the total value of loan after 4 years

Yn PMT No of periods Rate 13,000.00 4.00 5.00% 7 Value of loan ($56,031.63)

А в PMT 13000 No of periods 4 Rate 0.05 Value of loan =FV(C5, C4,C3)

Hence, total value of loan after 4 years = 56,031.6

This will be repaid over 10 years in monthly instalments

Total no of monthly instalments = 10*12 = 120

Monthly interest rate = (5%/12) = 0.42%

Using Excel's PMT function to arrive at monthly payments :-

A 12 13 14 15 PV Rate No of periods $56,031.63 0.42% 120.00 16 17 18 Monthly instalments ($594.30)

A PV Rate No of periods 56031.625 =5%/12 =10*12 17 18 Monthly instalments =PMT(C15,C16, C14) 19

Monthly instalments = $594.3